OpenStudy (anonymous):

A bookstore sold 18,000 paperbacks one month. This was 10% less than the number of paperbacks the store sold the previous month. The following equation represents this situation, where x represents the number of paperbacks sold the previous month. x - 0.1x = 18,000 How many paperbacks did the store sell in both months combined?

OpenStudy (aravindg):

:)

OpenStudy (anonymous):

Since 0.9x = 18,000, x = 18,000/0.9 = 180,000/9 = 20,000

OpenStudy (anonymous):

Where did you get 16,200 from?

OpenStudy (anonymous):

That answer is wrong. The correct answer for total books during both months is 18,000 + 20,000 = 38,000.

OpenStudy (anonymous):

sorry messed up.